OpenStudy (anonymous):

Find the interior angle sum for each polygon. Round your answer to the nearest tenth if necessary. One of the interior angle of a regular polygon is 165, find the number of sides of the polygon.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

use this formula: (n-2)*180/n , where n = number of sides

OpenStudy (campbell_st):

if the interior angle is 165 the exterior angle is 15 as they are supplementary the sum of the exterior angles is 360 so the number of sides(angles) = 360/15 the regular polygon has 24 sides

OpenStudy (anonymous):

165 = (n-2)*180/n n = 24

OpenStudy (campbell_st):

the interior angle sum is \[Angle Sum = (24 - 2)\times 180 \] the angle sum is 3960

OpenStudy (anonymous):

thanks guys! now one last question.. that's a little bit related. Find the number of sides of a regular polygon with interior angle sum of 2160.

OpenStudy (campbell_st):

ok... use the formula 2160 = (n -2) *180 \[\frac{2160}{180} = n - 2\] see if you can solve the above equation

OpenStudy (anonymous):

14? haah sorry I was gone for a long time :D

OpenStudy (anonymous):

uhm hello...?

ganeshie8 (ganeshie8):

correct :)